Please help, I have set this equation up several times and it is wrong! An athlete ran 6 mph and later cycled 20 mph to complete a 90 mile course. How far did she run and cycle
Answer by Alan3354(69443)

Please help, I have set this equation up several times and it is wrong! An athlete ran 6 mph and later cycled 20 mph to complete a 90 mile course in 6.25 hours. How far did she run and cycle
----------------------
d = r*t
t = d/r
d/6 + (90-d)/20 = 6.25
10d + 3(90-d) = 375
7d + 270 = 375
d = 15 miles running
75 miles cycling